OpenStudy (anonymous):

Help?????!! Given the function f(x) = x 2 - 2x, find f(-1) and f(1). Based on your answers, determine if the function is even. Explain your reasoning.

OpenStudy (anonymous):

y = x^2 - 2x solve for x and switch the variable y and x

OpenStudy (mertsj):

f(x) is even if f(x)=f(-x)

OpenStudy (mertsj):

So find f(1) Find f(-1) If the answers are the same, then the function is even.

OpenStudy (anonymous):

I dont know how to find them though..

OpenStudy (mertsj):

Replace x with 1. What do you get?

OpenStudy (anonymous):

f(x) and f(-x) ?

OpenStudy (mertsj):

y=2x. Find y if x = 1

OpenStudy (anonymous):

y=2

OpenStudy (mertsj):

y=x^2-2x Find y if x = 1

OpenStudy (anonymous):

y=-3 ?

OpenStudy (mertsj):

\[y=(1)^2-2(1)\]

OpenStudy (mertsj):

What is 1^2 ?

OpenStudy (anonymous):

1

OpenStudy (mertsj):

What is -2(1) ?

OpenStudy (anonymous):

So y =-1

OpenStudy (mertsj):

yes. f(x)=x^2-2x Find f(1) In other words...replace x with 1

OpenStudy (anonymous):

Oh okay I think I get it thank you!

OpenStudy (mertsj):

yw